OpenStudy (anonymous):

help please If f(x) = 3x + 1 and g(x) = 2x − 4, find f[g(2)].

OpenStudy (anonymous):

It's like substituting the function g(x) into the function f(x) \[f(x) = 3(2x-4) + 1\] \[f(x) = 6x-12 + 1\] \[f(x) = 6x - 11\]

OpenStudy (anonymous):

Substitute value of x=2 \[f(x) = 6(2) - 11\] \[f(x) = 12 - 11\] \[f(x) = 1\]

OpenStudy (anonymous):

thanks how about this one. If f(x) = −x − 1 and g(x) = 4x − 1, find f[g(−2)]

OpenStudy (anonymous):

\[f(x) = -x - 1\] \[f(x) = -(4x - 1) - 1\] \[f(x) = -4x + 1 - 1\] \[f(x) = -4x\] Substitute x=-2 \[f(x) = -4(-2)\] \[f(x) = 8\]